OpenStudy (anonymous):

Solve these equations by using elimination method: 2x + 9y = 4 3x + 7y = –7

OpenStudy (e.mccormick):

Do you know the meaning of the elimination method?

OpenStudy (anonymous):

Yes.

OpenStudy (e.mccormick):

OK, then where are you having trouble?

OpenStudy (anonymous):

Well I multiplied the first equation by -3 and the second by 2 and I got -6x - 27y = 4 6x + 14y = -14 - 13y = -10

OpenStudy (anonymous):

i need help with those things too !

OpenStudy (anonymous):

I don't think that this is going in the correct direction...Am I allowed to end up with a fraction?

OpenStudy (e.mccormick):

You forgot to multiply through the right side of the first equation.

OpenStudy (anonymous):

so how do you do it ?

OpenStudy (e.mccormick):

fractions are perfectly fine. You can solve with them, or have them as answers. However, in this case you just made a small math error.

OpenStudy (anonymous):

Ohhh. Ohkay. Thank you =)

OpenStudy (e.mccormick):

np. I am working out the fractional form just to show you.

OpenStudy (anonymous):

Ohkay! Thank you =)

OpenStudy (precal):